for the function v = 12 sin 40, calcualte the


1
Expert's answer
2022-03-24T07:50:13-0400

To solve the problem it is necessary, first, to determine the value of the Sin 40.

This can be done in the following ways:

1.   For example, in a right triangle

sin 40° = a / c, where

a - opposite leg;

c is the hypotenuse.

2.   You can just look at the sine table:

sin 40 ° = 0.6428 (up to 4 decimal places).

3.   You can count on the calculator.


We used the tables and checked with a calculator.

The following is the calculation result.

If the angle is determined in radians then the solution will be as follows:


"V=12*Sin(40Rad)=12*0.745=8.94".

 

If the angle is determined in degrees then the solution will be as follows:

 

"V=12*Sin(40Degree)=12*0.6428=7.7436".